OpenStudy (anonymous):

what is the solution of each proportion 7/9 = m/27

OpenStudy (gudden):

u mean after equating these... ??

OpenStudy (anonymous):

yes

OpenStudy (gudden):

is then 7/9 = m/27 cross multipling 7*27 =m*9 it means m = 7 * 3 = 21